OpenStudy (anonymous):

I need to use the shell method to find the volume of the of the enclosed portion in between these two curves. Pleas be elaborate! I keep getting 16 pi but apparently I need 8pi.

OpenStudy (anonymous):

y=x^2, y=6x-2x^2 I'm using r=x-0 and h=6x-2x^2

OpenStudy (anonymous):

Around which axis are you rotating this area?

OpenStudy (anonymous):

oh, sorry Y-axis

OpenStudy (phi):